# On balanced Hermitian threefolds with parallel Bismut torsion

### v1

@@ 3. $(3)$
For part (1), note that compact Chern flat manifolds are always balanced. The recent work [24] gives a satisfactory answer to (1). By the classic result of Boothby [10], any compact Chern flat manifold $(M^{n},g)$ is a quotient of a complex Lie group $G$ equipped with a left-invariant metric $\tilde{g}$ which is compatible with the complex structure of $G$, and $\tilde{g}$ is the lift of $g$. Theorem 1.2 in [24] states that, if $g$ is BTP, then $G$ must be reductive, and in fact it is the orthogonal direct product $G={\mathbb{C}}^{k}\times G_{1}\times\cdots\times G_{r}$ where each $G_{i}$ is a simple complex Lie group. Conversely, any reductive complex Lie group $G$ admits a left-invariant metric which is BTP (and is balanced and Chern flat).
For part (2), not much is known except a couple of partial results from [24]. Note that the flag threefold is $X=SU(3)/T^{2}$. Theorem 1.4 of [24] generalizes the Wallach threefold case to higher dimensions: for any $k\geq 2$, the metric $g$ naturally induced by the Killing form on full flag $X^{n}=SU(k+1)/T^{k}$ is (balanced, non-Kähler) BTP, and any non-Kähler BTP metric on $X^{n}$ is a constant multiple of $g$. Here $n=\frac{1}{2}k(k+1)$ is the complex dimension. Theorem 1.3 of [24] says the same thing holds if one replaces the type A full flag by any Kähler C-space of the form $Y=K/H$ where $K$ is a compact simple group and the isotropy representation of $H$ has at most two direct summands. Table 1 of [24] listed all such Kähler C-spaces. The Killing metric $g$ is Kähler when and only when $Y$ is a compact Hermitian symmetric space. For all others in Table 1, we get examples of non-Kähler, balanced BTP manifolds.
@@
To be more precise, let $(M^{3},g,J)$ be a primary balanced BTP threefold of middle type, then as we have seen before, it has a Vaisman companion $(M^{3},g,I)$, which is a particular type of Vaisman threefold. The specialty of $(M^{3},g,I)$ lies in the fact that its Bismut holonomy group $\mbox{Hol}^{b}(M)$ is $U(1)\!\times\!U(1)\!\times\!1$, or equivalently, the holomorphic tangent bundle $T_{M}$ of $(M^{3},g,I)$ is the orthogonal direct sum $L_{1}\oplus L_{2}\oplus L$ of complex line bundles and the decomposition is preserved by the Bismut connection $\nabla^{b}$. For a given (non-Kähler) Vaisman $n$-manifold, denote by $\eta$ the Gauduchon torsion $1$-form, which is the global $(1,0)$-form on $M^{n}$ defined by $\partial(\omega^{n-1})=-\eta\wedge\omega^{n-1}$. Denote by $\chi$ the vector field dual to $\eta$, namely, $\langle Y,\overline{\chi}\rangle=\eta(Y)$ for any type $(1,0)$ vector $Y$. It is easy to verify that $\chi$ is a holomorphic vector field of constant length, which also holds for general non-balanced BTP manifold as shown in [35, Proposition 1.7]. The above discussion motivates us to propose the following:
@@
Let $(M^{2m+1},g,I)$ be a compact Vaisman manifold of splitting type. Define an almost complex structure $J$ on $M$ by letting $J=I$ on $L_{1}\oplus L$ and $J=-I$ on $L_{2}$. Then $J$ is integrable, and the Hermitian manifold $(M^{2m+1},g,J)$ is balanced BTP. Furthermore, the Bismut connection of these two Hermitian manifolds coincide.
@@
The proof is analogous to the $3$-dimensional case, and we give it here for the sake of completeness. Write $n=2m+1$, and let $\nabla^{b}$ be the Bismut connection and $\chi$ be the holomorphic vector field on $(M^{n},g,I)$ dual to Gauduchon’s torsion $1$-form $\eta$. By [7], Vaisman manifolds are BTP, so $\nabla^{b}\chi=0$. Let $\{e_{1},\ldots,e_{n}\}$ be a local unitary frame of $(M^{n},g,I)$ so that $\lambda e_{n}=\chi$ where $\lambda=|\eta|>0$ is a global constant, $\{e_{1},\ldots,e_{m}\}$ spans $L_{1}$, and $\{e_{m\!+\!1},\ldots,e_{2m}\}$ spans $L_{2}$. Also let $\varphi$ be the coframe dual to $e$. For convenience, let us denote by $e^{\prime}$, $e^{\prime\prime}$ the column vector ($\,{}^{t}\!(e_{1},\ldots,e_{m})$ and ($\,{}^{t}\!(e_{m\!+\!1},\ldots,e_{2m})$, respectively, and similarly write $\varphi^{\prime}$, $\varphi^{\prime\prime}$ for the column vector ($\,{}^{t}\!(\varphi_{1},\ldots,\varphi_{m})$ and ($\,{}^{t}\!(\varphi_{m\!+\!1},\ldots,\varphi_{2m})$. Since both $L_{1}$ and $L_{2}$ are preserved by $\nabla^{b}$, we have $\nabla^{b}e^{\prime}=\theta^{b}_{1}e^{\prime}$, $\nabla^{b}e^{\prime\prime}=\theta^{b}_{1}e^{\prime\prime}$, and $\nabla^{b}e_{n}=0$ for some $m\times m$ matrices of $1$-forms $\theta^{b}_{1}$, $\theta^{b}_{2}$, so the Bismut connection matrix under $e$ is block-diagonal:
@@
where $\theta_{i}=\theta^{b}_{i}-\alpha I$ for $i=1$, $2$. By the definition of $J$ on $M^{n}$, it holds that $Jx=Ix$ for $x\in L_{1}\oplus L\oplus\overline{L}_{1}\oplus\overline{L}$ and $Jx=-Ix$ for $x\in L_{2}\oplus\overline{L}_{2}$. In order the verify that $J$ is integrable, we need to show that
(6.3) $[x,y]-[Jx,Jy]+J[Jx,y]+J[x,Jy]=0$
@@
(6.5) $02[e_{i},\overline{e}_{\alpha}]+2\sqrt{-1}\,J[e_{i},\overline{e}_{\alpha}]=$ 0 .
@@
$[e_{i},\overline{e}_{\alpha}]=\nabla^{c}_{e_{i}}\overline{e}_{\alpha}-\nabla^{c}_{\overline{e}_{\alpha}}e_{i}=\sum_{\beta=m+1}^{2m}\overline{\theta_{\alpha\beta}(\overline{e}_{i})}\,\overline{e}_{\beta}-\sum_{k=1}^{m}\theta_{ik}(\overline{e}_{\alpha})e_{k}$
@@
$[e_{n},\overline{e}_{\alpha}]=\sum_{\beta=m+1}^{2m}\overline{\theta_{\alpha\beta}(\overline{e}_{n})}\,\overline{e}_{\beta},$
as $\theta_{n\ast}$ are all $(1,0)$-forms. Therefore $[e_{i},\overline{e}_{\alpha}]$ is always a linear combination of $\overline{e}_{\beta}$ and $e_{k}$, for $1\leq k\leq m$ and $m\!+\!1\leq\beta\leq 2m$. Since $J\overline{e}_{\beta}=-I\overline{e}_{\beta}=\sqrt{-1}\overline{e}_{\beta}$ and $Je_{k}=Ie_{k}=\sqrt{-1}e_{k}$, we have $J[e_{i},\overline{e}_{\alpha}]=\sqrt{-1}\,[e_{i},\overline{e}_{\alpha}]$, hence (6.5) holds. This completes the proof that $J$ is integrable.
Next let us show that the Bismut connection ^$\hat{\nabla}^{b}$ of the Hermitian manifold $(M^{n},g,J)$ will coincide with the Bismut connection $\nabla^{b}$ of the original Vaisman manifold $(M^{n},g,I)$. Let $\varepsilon_{i}=e_{i}$ and $\psi_{i}=\varphi_{i}$ for $1\leq i\leq m$ or $i=n$, while $\varepsilon_{\alpha}=\overline{e}_{\alpha}$ and $\psi_{\alpha}=\overline{\varphi}_{\alpha}$ for $m\!+\!1\leq\alpha\leq 2m$. Then $\varepsilon$ becomes a unitary frame for $(M^{n},g,J)$ and $\psi$ is its dual coframe. Again write $\psi^{\prime}=\,^{t}\!(\psi_{1},\ldots,\psi_{m})$ and $\psi^{\prime\prime}=\,^{t}\!(\psi_{m\!+\!1},\ldots,\psi_{2m})$ for the column vectors. By taking the complex conjugation in the middle portion of (6.2), we obtain
@@
Hence ^$\hat{\nabla}^{b}e^{\prime}=\hat{\nabla}^{b}\varepsilon^{\prime}=\hat{\theta}^{b}_{1}\varepsilon^{\prime}=\theta^{b}_{1}\varepsilon^{\prime}=\theta^{b}_{1}e^{\prime}=\nabla^{b}e^{\prime}$, ^$\hat{\nabla}^{b}e_{n}=\hat{\nabla}^{b}\varepsilon_{n}=0=\nabla^{b}e_{n}$, and
^$\hat{\nabla}^{b}e^{\prime\prime}=\hat{\nabla}^{b}\overline{\varepsilon^{\prime\prime}}=\overline{\hat{\theta}^{b}_{2}\varepsilon^{\prime\prime}}=\overline{\overline{\theta^{b}_{2}}\varepsilon^{\prime\prime}}=\theta^{b}_{2}e^{\prime\prime}=\nabla^{b}e^{\prime\prime}.$
Therefore, ^$\hat{\nabla}^{b}=\nabla^{b}$, namely, Hermitian manifolds $(M^{n},g,I)$ and $(M^{n},g,J)$ share the same Bismut connection, thus the latter is a balanced BTP manifold. This completes the proof of the theorem. ∎
@@ 7. Appendix
By [35, Proposition 2.5], we know that $\nabla^{b}_{1,0}T=0$ implies that the right hand side of the equality above is zero. Therefore $\nabla^{b}_{1,0}T=0$ indicates $\nabla^{c}_{1,0}T=0$. It implies if $g$ is BTP, then the $(1,0)$-part of Chern covariant differentiation of the torsion vanishes, that is, $T^{j}_{ik;\ell}=0$, where the index after the semicolon stands for covariant derivative with respect to the Chern connection. The $(0,1)$-part of Chern covariant differentiation of torsion, on the other hand, is given by
@@
Denote by $R$ the Riemannian curvature tensor, namely, the curvature of the Levi-Civita connection of $g$. Note that the symbol $T_{ik}^{j}$ defined at the beginning of §2 is two times of that in [33]. So by [33, Lemma 7] and the equality above, we obtain
@@
Then let us specialize to the Wallach threefold $(X,g)$ at the origin 00. Recall from §4 that $g_{i\overline{j}}=\delta_{ij}$ at the origin and all the components of $T$ vanish except $T^{2}_{13}=1$, so the right hand side of (7.1) is zero, hence $R_{ijk\overline{\ell}}=0$. By the formulae on $R^{c}_{i\overline{j}k\overline{\ell}}$ obtained in §4, we get
@@
where we recall that $R_{i\overline{j}k\overline{\ell}}=R_{k\overline{\ell}i\overline{j}}$ always holds for the Riemannian curvature $R$. Now we compute the sectional curvature of $R$. Let $x$, $y$ be any two real tangent vector of $X$ at the origin 00 satisfying $x\wedge y\neq 0$. Write $x=X+\overline{X}$ and $y=Y+\overline{Y}$ for type $(1,0)$ tangent vectors $X$ and $Y$. We have
@@
$-R_{XY\overline{X}\overline{Y}}=-R_{X\overline{X}Y\overline{Y}}+R_{Y\overline{X}X\overline{Y}}.$
@@
$=$ $0\displaystyle 2\{\mbox{Im}(X_{i}\overline{Y}_{i})-\mbox{Im}(X_{k}\overline{Y}_{k})\}^{2}+\frac{1}{4}|X_{i}Y_{k}-X_{k}Y_{i}|^{2}\ \geq$ 0 .
Similarly, for $(ik)=(12)$ or $(23)$, $12a_{ik}-b_{ik}=$ 1, $a_{ik}+b_{ik}=\frac{5}{4}$, so
@@
$=$ $0\displaystyle 2\{\mbox{Im}(X_{i}\overline{Y}_{i})+\mbox{Im}(X_{k}\overline{Y}_{k})\}^{2}+\frac{1}{4}|X_{i}\overline{Y}_{k}-Y_{i}\overline{X}_{k}|^{2}\ \geq$ 0 .

### v2

@@ 3. $(3)$
For part (1), note that compact Chern flat manifolds are always balanced. The recent work [PZ25] gives a satisfactory answer to (1). By the classic result of Boothby [Boo58], any compact Chern flat manifold $(M^{n},g)$ is a quotient of a complex Lie group $G$ equipped with a left-invariant metric $\tilde{g}$ which is compatible with the complex structure of $G$, and $\tilde{g}$ is the lift of $g$. Theorem 1.2 in [PZ25] states that, if $g$ is BTP, then $G$ must be reductive, and in fact it is the orthogonal direct product $G={\mathbb{C}}^{k}\times G_{1}\times\cdots\times G_{r}$ where each $G_{i}$ is a simple complex Lie group. Conversely, any reductive complex Lie group $G$ admits a left-invariant metric which is BTP (and is balanced and Chern flat).
For part (2), not much is known except a couple of partial results from [PZ25]. Note that the flag threefold is $X=SU(3)/T^{2}$. Theorem 1.4 of [PZ25] generalizes the Wallach threefold case to higher dimensions: for any $k\geq 2$, the metric $g$ naturally induced by the Killing form on full flag $X^{n}=SU(k+1)/T^{k}$ is (balanced, non-Kähler) BTP, and any non-Kähler BTP metric on $X^{n}$ is a constant multiple of $g$. Here $n=\frac{1}{2}k(k+1)$ is the complex dimension. Theorem 1.3 of [PZ25] says the same thing holds if one replaces the type A full flag by any Kähler C-space of the form $Y=K/H$ where $K$ is a compact simple group and the isotropy representation of $H$ has at most two direct summands. Table 1 of [PZ25] listed all such Kähler C-spaces. The Killing metric $g$ is Kähler when and only when $Y$ is a compact Hermitian symmetric space. For all others in Table 1, we get examples of non-Kähler, balanced BTP manifolds.
@@
To be more precise, let $(M^{3},g,J)$ be a primary balanced BTP threefold of middle type, then as we have seen before, it has a Vaisman companion $(M^{3},g,I)$, which is a particular type of Vaisman threefold. The specialty of $(M^{3},g,I)$ lies in the fact that its Bismut holonomy group $\mbox{Hol}^{b}(M)$ is $U(1)\!\times\!U(1)\!\times\!1$, or equivalently, the holomorphic tangent bundle $T_{M}$ of $(M^{3},g,I)$ is the orthogonal direct sum $L_{1}\oplus L_{2}\oplus L$ of complex line bundles and the decomposition is preserved by the Bismut connection $\nabla^{b}$. For a given (non-Kähler) Vaisman $n$-manifold, denote by $\eta$ the Gauduchon torsion $1$-form, which is the global $(1,0)$-form on $M^{n}$ defined by $\partial(\omega^{n-1})=-\eta\wedge\omega^{n-1}$. Denote by $\chi$ the vector field dual to $\eta$, namely, $\langle Y,\overline{\chi}\rangle=\eta(Y)$ for any type $(1,0)$ vector $Y$. It is easy to verify that $\chi$ is a holomorphic vector field of constant length, which also holds for general non-balanced BTP manifold as shown in [ZZ24+, Proposition 1.7]. The above discussion motivates us to propose the following:
@@
Let $(M^{2m+1},g,I)$ be a compact Vaisman manifold of splitting type. Define an almost complex structure $J$ on $M$ by letting $J=I$ on $L_{1}\oplus L$ and J=IJ=-I on $L_{2}$. Then $J$ is integrable, and the Hermitian manifold $(M^{2m+1},g,J)$ is balanced BTP. Furthermore, the Bismut connection of these two Hermitian manifolds coincide.
@@
[…]
@@
where $\theta_{i}=\theta^{b}_{i}-\alpha I$ for $i=1$, 22. By the definition of $J$ on $M^{n}$, it holds that $Jx=Ix$ for $x\in L_{1}\oplus L\oplus\overline{L}_{1}\oplus\overline{L}$ and Jx=IxJx=-Ix for $x\in L_{2}\oplus\overline{L}_{2}$. In order the verify that $J$ is integrable, we need to show that
(6.3) [x,y][Jx,Jy]+J[Jx,y]+J[x,Jy]=0[x,y]-[Jx,Jy]+J[Jx,y]+J[x,Jy]=0
@@
(6.5) $2[e_{i},\overline{e}_{\alpha}]+2\sqrt{-1}\,J[e_{i},\overline{e}_{\alpha}]=0.$
@@
$e_{i},\overline{e}_{\alpha}]=\nabla^{c}_{e_{i}}\overline{e}_{\alpha}-\nabla^{c}_{\overline{e}_{\alpha}}e_{i}=\sum_{\beta=m+1}^{2m}\overline{\theta_{\alpha\beta}(\overline{e}_{i})}\,\overline{e}_{\beta}-\sum_{k=1}^{m}\theta_{ik}(\overline{e}_{\alpha})e_{k}$
@@
$e_{n},\overline{e}_{\alpha}]=\sum_{\beta=m+1}^{2m}\overline{\theta_{\alpha\beta}(\overline{e}_{n})}\,\overline{e}_{\beta},$
as $\theta_{n\ast}$ are all $(1,0)$-forms. Therefore $e_{i},\overline{e}_{\alpha}]$ is always a linear combination of $\overline{e}_{\beta}$ and $e_{k}$, for $1\leq k\leq m$ and $m\!+\!1\leq\beta\leq 2m$. Since $J\overline{e}_{\beta}=-I\overline{e}_{\beta}=\sqrt{-1}\overline{e}_{\beta}$ and $Je_{k}=Ie_{k}=\sqrt{-1}e_{k}$, we have $J[e_{i},\overline{e}_{\alpha}]=\sqrt{-1}\,[e_{i},\overline{e}_{\alpha}]$, hence (6.5) holds. This completes the proof that $J$ is integrable.
Next let us show that the Bismut connection $\hat{\nabla}^{b}$ of the Hermitian manifold $(M^{n},g,J)$ will coincide with the Bismut connection $\nabla^{b}$ of the original Vaisman manifold $(M^{n},g,I)$. Let $\varepsilon_{i}=e_{i}$ and $\psi_{i}=\varphi_{i}$ for $1\leq i\leq m$ or $i=n$, while $\varepsilon_{\alpha}=\overline{e}_{\alpha}$ and $\psi_{\alpha}=\overline{\varphi}_{\alpha}$ for $m\!+\!1\leq\alpha\leq 2m$. Then $\varepsilon$ becomes a unitary frame for $(M^{n},g,J)$ and $\psi$ is its dual coframe. Again write $\psi^{\prime}=\,^{t}\!(\psi_{1},\ldots,\psi_{m})$ and $\psi^{\prime\prime}=\,^{t}\!(\psi_{m\!+\!1},\ldots,\psi_{2m})$ for the column vectors. By taking the complex conjugation in the middle portion of (6.2), we obtain
@@
Hence $\hat{\nabla}^{b}e^{\prime}=\hat{\nabla}^{b}\varepsilon^{\prime}=\hat{\theta}^{b}_{1}\varepsilon^{\prime}=\theta^{b}_{1}\varepsilon^{\prime}=\theta^{b}_{1}e^{\prime}=\nabla^{b}e^{\prime}$, $\hat{\nabla}^{b}e_{n}=\hat{\nabla}^{b}\varepsilon_{n}=0=\nabla^{b}e_{n}$, and
$\hat{\nabla}^{b}e^{\prime\prime}=\hat{\nabla}^{b}\overline{\varepsilon^{\prime\prime}}=\overline{\hat{\theta}^{b}_{2}\varepsilon^{\prime\prime}}=\overline{\overline{\theta^{b}_{2}}\varepsilon^{\prime\prime}}=\theta^{b}_{2}e^{\prime\prime}=\nabla^{b}e^{\prime\prime}.$
Therefore, $\hat{\nabla}^{b}=\nabla^{b}$, namely, Hermitian manifolds $(M^{n},g,I)$ and $(M^{n},g,J)$ share the same Bismut connection, thus the latter is a balanced BTP manifold. This completes the proof of the theorem. ∎
@@ 7. Appendix
By [ZZ24+, Proposition 2.5], we know that $\nabla^{b}_{1,0}T=0$ implies that the right hand side of the equality above is zero. Therefore $\nabla^{b}_{1,0}T=0$ indicates $\nabla^{c}_{1,0}T=0$. It implies if $g$ is BTP, then the $(1,0)$-part of Chern covariant differentiation of the torsion vanishes, that is, $T^{j}_{ik;\ell}=0$, where the index after the semicolon stands for covariant derivative with respect to the Chern connection. The $(0,1)$-part of Chern covariant differentiation of torsion, on the other hand, is given by
@@
Denote by $R$ the Riemannian curvature tensor, namely, the curvature of the Levi-Civita connection of $g$. Note that the symbol $T_{ik}^{j}$ defined at the beginning of §2 is two times of that in [YZ18]. So by [YZ18, Lemma 7] and the equality above, we obtain
@@
Then let us specialize to the Wallach threefold $(X,g)$ at the origin 0. Recall from §4 that $g_{i\overline{j}}=\delta_{ij}$ at the origin and all the components of $T$ vanish except $T^{2}_{13}=1$, so the right hand side of (7.1) is zero, hence $R_{ijk\overline{\ell}}=0$. By the formulae on $R^{c}_{i\overline{j}k\overline{\ell}}$ obtained in §4, we get
@@
where we recall that $R_{i\overline{j}k\overline{\ell}}=R_{k\overline{\ell}i\overline{j}}$ always holds for the Riemannian curvature $R$. Now we compute the sectional curvature of $R$. Let $x$, $y$ be any two real tangent vector of $X$ at the origin 0 satisfying $x\wedge y\neq 0$. Write $x=X+\overline{X}$ and $y=Y+\overline{Y}$ for type $(1,0)$ tangent vectors $X$ and $Y$. We have
@@
$R_{XY\overline{X}\overline{Y}}=-R_{X\overline{X}Y\overline{Y}}+R_{Y\overline{X}X\overline{Y}}.$
@@
$=$ $2\{\mbox{Im}(X_{i}\overline{Y}_{i})-\mbox{Im}(X_{k}\overline{Y}_{k})\}^{2}+\frac{1}{4}|X_{i}Y_{k}-X_{k}Y_{i}|^{2}\ \geq\ 0.$
Similarly, for $(ik)=(12)$ or $(23)$, $2a_{ik}-b_{ik}=1$, $a_{ik}+b_{ik}=\frac{5}{4}$, so
@@
$=$ $2\{\mbox{Im}(X_{i}\overline{Y}_{i})+\mbox{Im}(X_{k}\overline{Y}_{k})\}^{2}+\frac{1}{4}|X_{i}\overline{Y}_{k}-Y_{i}\overline{X}_{k}|^{2}\ \geq\ 0.$